Hi tutors;
I need some help with proportions. I have a problem requiring me to find a proportion in feet, while the other side is in inches. This is the proportion:
8 inches/20 inches = 22 ft/x ft
I am confused on this one. :/

Answer by rfer(16322)

proportions is one area you don't need to be concerned about same values
8/20=22/x
8x=440
x=55
----------
8/20=22*12/x
8x=5280 inches
x=660 inches
660/12=55 ft